# LoRA-IR: Taming Low-Rank Experts for Efficient All-in-One Image Restoration

### v1

@@ Abstract
Prompt-based all-in-one image restoration (IR) frameworks have achieved remarkable performance by incorporating degradation-specific information into prompt modules. Nevertheless, handling the complex and diverse degradations encountered in real-world scenarios remains a significant challenge. To address this challenge, we propose LoRA-IR, a flexible framework that dynamically leverages compact low-rank experts to facilitate efficient all-in-one image restoration. Specifically, LoRA-IR consists of two training stages: degradation-guided pre-training and parameter-efficient fine-tuning. In the pre-training stage, we enhance the pre-trained CLIP model by introducing a simple mechanism that scales it to higher resolutions, allowing us to extract robust degradation representations that adaptively guide the IR network. In the fine-tuning stage, we refine the pre-trained IR network using low-rank adaptation (LoRA). Built upon a Mixture-of-Experts (MoE) architecture, LoRA-IR dynamically integrates multiple low-rank restoration experts through a degradation-guided router. This dynamic integration mechanism significantly enhances our model’s adaptability to diverse and unknown degradations in complex real-world scenarios. Extensive experiments demonstrate that LoRA-IR achieves state-of-the-art performance across 14 image restoration tasks and 29 benchmarks. Code and pre-trained models will be available at: https://github.com/shallowdream204/LoRA-IR.
@@ 1 Introduction
Image restoration (IR) is a fundamental task in computer vision, aiming to recover high-quality (HQ) images from degraded low-quality (LQ) inputs. In recent years, significant progress has been achieved with specialized restoration networks targeting specific degradations such as adverse weather removal [15, 7], denoising [83], deblurring [58], and low-light enhancement [70]. However, in practical applications like autonomous driving and outdoor surveillance [40, 96], images are often simultaneously affected by multiple complex degradations, including haze, rain, snow, low-light conditions, and motion blur. These complex degradations not only degrade image quality but also severely impair the performance of downstream vision tasks, posing significant challenges to the safety and reliability of systems. Existing specialized models designed for single-task restoration struggle to generalize effectively in such unpredictable and variable environments.
@@
Figure 1: PSNR comparison with state-of-the-art all-in-one methods across 8 tasks (Tab. 4 and Tab. 9).
@@
Figure 2: Conceptual comparison of all-in-one frameworks. (a) Multi-Encoder Structures: Use multiple encoders to extract features, but redundancy reduces computational efficiency. (b) Prompt-Based Methods: Employ lightweight prompts for degradation-specific features, improving efficiency. However, static network structures limit their ability to handle unknown complex degradations. (c) Our Proposed Framework: Dynamically and sparsely combines low-rank restoration experts. This design maintains computational efficiency while adapting flexibly to diverse degradation types, enhancing real-world performance.
@@
To overcome the limitations of specialized models, there is increasing interest in developing all-in-one frameworks capable of handling diverse degradations. Early approaches, such as multi-encoder architectures [24] (Fig. 2 (a)), employ separate encoders for different degradation types. While effective in handling multiple degradations, their redundant structures lead to a large number of parameters, hindering scalability and efficiency. More recent methods adopt prompt-based frameworks [48, 33, 36, 2] (Fig. 2 (b)), encoding degradation-specific information into lightweight prompts to guide a shared network. However, relying solely on lightweight prompts and a static shared network may not fully capture the fine-grained details and specific patterns associated with different degradations, leading to suboptimal restoration results. Furthermore, potential correlations and shared features among different degradations—such as common patterns in adverse weather conditions [95, 80]—are not extensively leveraged. Exploring these correlations could be key to enhancing model adaptability and effectiveness in complex real-world scenarios.
@@
In this work, we propose LoRA-IR, a flexible framework for efficient all-in-one image restoration (Fig. 2 (c)). Motivated by the success of Low-Rank Adaptation (LoRA) [14] in parameter-efficient fine-tuning, this paper explores the use of diverse low-rank experts to model degradation characteristics and correlations efficiently. LoRA-IR consists of two stages of training, both guided by a proposed degradation-guided router (DG-Router). DG-Router is based on the powerful vision-language model CLIP [51], which has demonstrated strong representational capabilities across a wide range of high-level vision tasks [29, 90]. However, when applied to low-level tasks, its limited input resolution inevitably leads to suboptimal performance when handling high-resolution LQ images. To this end, we introduce a simple yet effective method for scaling CLIP to high resolution. Our approach involves downsampling the image and applying a sliding window technique to capture both global and local detail representations, which are subsequently fused using lightweight MLPs. With minimal training parameters and a short training time, DG-Router provides robust degradation representations and probabilistic guidance for the training of LoRA-IR.
@@
In the first stage, we use the degradation representations provided by the DG-Router to guide the pre-training of the IR network. The degradation representations dynamically modulate features within the IR network through the proposed degradation-guided adaptive modulator (DAM). In the second stage, we fine-tune the IR network obtained from the first stage using LoRA. Based on the Mixture-of-Expert (MoE) [53] structure, we construct a set of low-rank restoration experts. Leveraging the probabilistic guidance of the DG-Router, we sparsely select different LoRA experts to dynamically adjust the IR network. Different experts enhance the network’s ability to capture degradation-specific knowledge, while their collaboration equips the network with the capability to learn correlations between various degradations. The dynamic network structure enables LoRA-IR to adapt to diverse degradations and improves its generalization capabilities. As shown in Fig. 1, LoRA-IR outperforms all compared state-of-the-art all-in-one methods and demonstrates favorable generalizability in handling complex real-world scenarios.
@@
  • We propose LoRA-IR, a simple yet powerful baseline for all-in-one restoration. LoRA-IR employs a novel mixture of low-rank expert structures, boosting architectural flexibility while maintaining computational efficiency.
@@ 2 Related Work
2.1 Image Restoration
Image restoration for known degradations has been extensively studied [27, 77, 78, 65, 4, 82, 26, 71, 9]. Recently, there has been significant interest in all-in-one frameworks within the community [81, 95, 60]. AirNet [22] is the pioneering work in all-in-one IR, utilizing contrastive learning to capture degradation information. Recent SOTA methods are mostly based on prompt learning, using lightweight prompts to encode degradation information. PromptIR [48] proposes a plug-and-play prompt module to guide the restoration process. DA-CLIP [36] utilizes a prompt learning module to incorporate degradation embeddings. MPerceiver [2] introduces a multi-modal prompt learning approach to harness Stable Diffusion priors. Despite achieving promising results, most existing methods use fixed network architectures, which may limit their ability to cover complex real-world scenarios.
2.2 Vision-Language Models
Vision-language models (VLMs) have demonstrated strong performance across various multi-modal and vision-only tasks [51, 8, 29]. Among them, CLIP, as a powerful VLM, has demonstrated impressive zero-shot and few-shot capabilities across various high-level vision tasks [67, 93, 66]. However, in low-level vision tasks, CLIP’s capabilities have been relatively less explored. DA-CLIP [36] is the first to introduce CLIP into all-in-one IR. It adopts a ControlNet-style [86] structure, using contrastive learning with image-text pairs to fine-tune CLIP. In this work, we focus on exploring CLIP’s visual representation capabilities to achieve efficient degradation representation. Compared to DA-CLIP (Tab. 7), our proposed DG-Router requires 64×64$\times$64 × fewer learning parameters and 4×4$\times$4 × less training time, while achieving superior performance.
2.3 Parameter-efficient Fine-tuning
With the rise of large foundational models [51, 18, 1] in modern deep learning, the community has increasingly shifted its focus towards parameter-efficient fine-tuning (PEFT) methods for effective model adaptation. Among these, prompt learning [20, 90] and Low-Rank Adaptation (LoRA) [14] are two prominent and widely used PEFT methods. As discussed above, prompt learning has been widely applied in low-level vision tasks. LoRA posits that the weight changes during model adaptation follow a low-rank structure and incorporates trainable rank decomposition matrices into the pre-trained model. Specifically, the change matrix is re-parameterized into the product of two low-rank matrices: $W=W_{0}+\Delta W=W_{0}+sBA,$ where $W_{0}$ represents pre-trained weight matrix, $B\in\mathbb{R}^{m\times r}$ and $A\in\mathbb{R}^{r\times n}$ are low-rank matrices, $s=\frac{\alpha}{r}$ is the scaling factor. In this work, we first introduce LoRA into the all-in-one frameworks to facilitate efficient image restoration.
@@ 3 Method
As shown in Fig. 4, the image restoration network is based on the commonly used U-Net [65, 78, 4] structure, consisting of stacked encoder, middle, and decoder blocks. Our LoRA-IR consists of two training stages: degradation-guided pre-training and parameter-efficient fine-tuning, both directed by the proposed degradation-guided router (DG-Router). We first introduce the CLIP-based DG-Router in Sec. 3.1, which is used to extract robust degradation representations and provide probabilistic estimates to guide the training of LoRA-IR. Then we detail the pre-training process of LoRA-IR in Sec. 3.2. Finally, we describe the fine-tuning process in Sec. 3.3.
@@ 3.1 Degradation-guided Router
As shown in Fig. 4 (a), DG-Router uses a pre-trained CLIP image encoder to extract rich features from LQ images. The pre-trained CLIP image encoder typically limits input images to a smaller resolution (e.g., $224\times 224$). When handling higher-resolution images, a common approach [36] is to downsample the image to the resolution supported by CLIP using the processor. While this may have minimal impact on perception-based high-level classification tasks, significant downsampling can potentially lead to the loss of critical degradation information in pixel-level regression tasks like image restoration. Fig. 3 illustrates the results after the CLIP processor has processed the LQ images. It can be seen that significant downsampling leads to a substantial loss of degradation information in the LQ images, making it challenging to effectively extract degradation representations from the CLIP output features.
To address this issue, we propose a simple yet effective mechanism for scaling up the input resolution. For input LQ image $I_{LQ}\in\mathbb{R}^{H\times W\times 3}$, we use a sliding window to partition the image into small local patches $I_{slide}\in\mathbb{R}^{M\times H_{c}\times W_{c}\times 3}$, where $M$ is the number of patches, $H_{c}\times W_{c}$ is the CLIP supported resolution. Both $I_{slide}$ and down-sampled image $I_{down}\in\mathbb{R}^{H_{c}\times W_{c}\times 3}$ are fed into the image encoder simultaneously, obtaining output features $e^{slide}\in\mathbb{R}^{M\times C_{clip}}$ and $e^{down}\in\mathbb{R}^{C_{clip}}$. As depicted in Fig. 4 (a), after pooling $e^{slide}$, we concatenate them and pass them into a two-layer MLP to obtain the CLIP extracted degradation embedding $e^{clip}$, which can be formulated as
@@
After feeding $e^{clip}$ into the classification head, we ultimately obtain the degradation prediction probability $w\in\mathbb{R}^{n}$, where $n$ is the number of degradation types. Without bells and whistles, the DG-Router is optimized using standard cross-entropy loss, with the only parameters being the classification head and the two-layer MLP. Once training is complete, all parameters of the DG-Router are frozen and no longer updated.
@@ 3.2 Degradation-guided Pre-training
where $\odot$ denotes the element-wise multiplication, $\mathrm{MLP_{shared}}$ denotes the MLP projector shared across different blocks, $\mathrm{LN}$ denotes LayerNorm, $x_{in}$ is the original feature in the IR network, and $x_{out}$ is the feature after modulation. Through the modulation of DAM, the robust degradation representations from the DG-Router can effectively enhance the degradation-specific knowledge of the IR network during pre-training.
3.3 Parameter-efficient Fine-tuning
In the fine-tuning stage, we aim to utilize the Low-Rank Adaptation (LoRA) technique to model degradation characteristics and correlations efficiently, enhancing the model’s ability to handle training-seen degradations and improving its robustness for real-world training-unseen degradations.
As shown in Fig. 4 (c), built upon the Mixture-of-Expert (MoE) architecture, we establish a collection composed of low-rank restoration experts. We have a total of $n$ low-rank experts $\{E_{1},E_{2},\cdots,E_{n}\}$, each expert is a learnable lightweight LoRA weight of the pre-trained restoration network from the first stage, specialized in handling a specific type of degradation.
For a given input LQ image, DG-Router can predict its degradation probability $w\in\mathbb{R}^{n}$. $w$ is used as the score of different experts to select the appropriate expert for the restoration process. We sparsely select the top-$k$ highest-scoring experts as the most relevant ones and achieve the final restoration result through their dynamic collaboration, which can be formulated as
where $PreMod$ denotes the pre-trained module in stage1, $\varphi(i)$ denotes the index of the $i$-th selected expert, $w^{\prime}\in\mathbb{R}^{n}$ represents the result of reapplying softmax normalization to the scores of the selected top-$k$ experts (the weights of the experts that were not selected are set to 00).
Note that the sparse selection mechanism in Eq. (3) grants LoRA-IR a flexible network structure, enhancing its ability to represent degradation-specific knowledge. The dynamic combination mechanism, on the other hand, facilitates collaboration among different restoration experts, effectively modeling the commonalities and correlations between various degradations. The design of the low-rank experts ensures the efficiency of LoRA-IR, enabling it to achieve high-performance all-in-one image restoration in a computationally efficient manner.
4 Experiments
Table 1: [Setting I] Quantitative comparisons for 4-task adverse weather removal. The top halves of the tables show task-specific results, while the bottom halves present all-in-one evaluations. LoRA-IR surpasses recent SOTA all-in-one techniques, including MPerceiver [2] and Histoformer [56], across all evaluated datasets and metrics.
@@ 4 Experiments
LoRA-IR (Ours) 33.39 0.9489

### v2

@@ Abstract
Prompt-based all-in-one image restoration (IR) frameworks have achieved remarkable performance by incorporating degradation-specific information into prompt modules. Nevertheless, handling the complex and diverse degradations encountered in real-world scenarios remains a significant challenge. To tackle this, we propose LoRA-IR, a flexible framework that dynamically leverages compact low-rank experts to facilitate efficient all-in-one image restoration. Specifically, LoRA-IR consists of two training stages: degradation-guided pre-training and parameter-efficient fine-tuning. In the pre-training stage, we enhance the pre-trained CLIP model by introducing a simple mechanism that scales it to higher resolutions, allowing us to extract robust degradation representations that adaptively guide the IR network. In the fine-tuning stage, we refine the pre-trained IR network through low-rank adaptation (LoRA). Built upon a Mixture-of-Experts (MoE) architecture, LoRA-IR dynamically integrates multiple low-rank restoration experts through a degradation-guided router. This dynamic integration mechanism significantly enhances our model’s adaptability to diverse and unknown degradations in complex real-world scenarios. Extensive experiments demonstrate that LoRA-IR achieves SOTA performance across 14 IR tasks and 29 benchmarks, while maintaining computational efficiency. Code and pre-trained models will be available at: https://github.com/shallowdream204/LoRA-IR.
@@ 1 Introduction
Figure 1: PSNR comparison with state-of-the-art all-in-one methods across 8 image restoration tasks (Tab. 4 and Tab. 9).
@@
Figure 2: Conceptual comparison of all-in-one frameworks. (a) Multi-Encoder Structures: Use multiple encoders to extract features, but redundancy reduces model efficiency. (b) Prompt-Based Methods: Employ lightweight prompts for degradation-specific features, improving efficiency. However, static network structures limit their ability to handle unknown complex degradations. (c) Our Proposed Framework: Self-adaptively and sparsely combines low-rank restoration experts. This design preserves model efficiency while enabling self-adaptation to various degradation types, thereby enhancing its real-world performance.
Image restoration (IR) is a fundamental task in computer vision, aiming to recover high-quality (HQ) images from degraded low-quality (LQ) inputs. In recent years, significant progress has been achieved with specialized restoration networks targeting specific degradations [7, 83, 58, 70]. However, in practical applications like autonomous driving and outdoor surveillance [40, 96], images are often simultaneously affected by multiple complex degradations, including haze, rain, low-light conditions, motion blur, etc. These intricate degradations not only degrade image quality but also severely impair the performance of downstream vision tasks, posing significant challenges to the safety and reliability of such systems. Specialized models designed for single-task restoration often struggle to generalize effectively in these unpredictable and variable environments.
@@
To overcome the limitations of specialized IR models, there is growing interest in developing all-in-one frameworks capable of handling diverse degradations. Early approaches, such as multi-encoder architectures [24] (Fig. 2 (a)), employ separate encoders for different degradation types. While effective in handling multiple degradations, their redundant structures lead to a large number of parameters, hindering scalability and efficiency. More recent state-of-the-art methods adopt prompt-based frameworks [48, 33, 36, 2] (Fig. 2 (b)), encoding degradation-specific information into lightweight prompts to guide a shared network. However, relying solely on lightweight prompts and a static shared network may not fully capture the fine-grained details and specific patterns associated with different degradations, leading to suboptimal restoration results. Furthermore, potential correlations and shared features among different degradations—such as common patterns in adverse weather conditions [95, 80]—are not extensively leveraged. Leveraging these correlations could be the key to enhancing model adaptability and effectiveness in complex real-world scenarios.
@@
In this work, we propose LoRA-IR, a flexible framework for efficient all-in-one image restoration (Fig. 2 (c)). Motivated by the success of Low-Rank Adaptation (LoRA) [15] in parameter-efficient fine-tuning, we explore the use of diverse low-rank experts to model degradation characteristics and correlations efficiently. LoRA-IR involves two training stages, both guided by the proposed Degradation-Guided Router (DG-Router). DG-Router is based on the powerful vision-language model CLIP [51], which has demonstrated strong representation capabilities across a wide range of high-level vision tasks [29, 90]. However, when applied to low-level tasks, its limited input resolution inevitably leads to suboptimal performance when handling high-resolution LQ images. To this end, we introduce a simple yet effective method for scaling CLIP to high resolution. Our approach involves downsampling the image and applying a sliding window technique to capture both global and local detail representations, which are subsequently fused using lightweight MLPs. With minimal trainable parameters and a short training time, DG-Router provides robust degradation representations and probabilistic guidance for the training of LoRA-IR.
@@
In the first stage, we use the degradation representations provided by DG-Router to guide the pre-training of the IR network. The degradation representations dynamically modulate features within the IR network through the proposed Degradation-guided Adaptive Modulator (DAM). In the second stage, we fine-tune the IR network obtained from the first stage using LoRA. Based on the Mixture-of-Experts (MoE) [53] structure, we construct a set of low-rank restoration experts. Leveraging the probabilistic guidance of the DG-Router, we sparsely select different LoRA experts to adaptively adjust the IR network. Each expert enhances the network’s ability to capture degradation-specific knowledge, while their collaboration equips the network with the capability to learn correlations between various degradations. The self-adaptive network structure enables LoRA-IR to adapt to diverse degradations and improves its generalization capabilities. As shown in Fig. 1, LoRA-IR outperforms all compared state-of-the-art all-in-one methods and demonstrates favorable generalizability in handling complex real-world scenarios.
@@
  • We propose LoRA-IR, a simple yet effective baseline for all-in-one IR. LoRA-IR leverages a novel mixture of low-rank experts structure, enhancing architectural adaptability while maintaining computational efficiency.
@@ 2 Related Work
Image Restoration. Image restoration for known degradations has been extensively studied [27, 77, 78, 65, 4, 82, 26, 71, 10]. Recently, there has been significant interest in all-in-one frameworks within the community [81, 95, 60, 9]. AirNet [22] is the pioneering work in all-in-one IR, utilizing contrastive learning to capture degradation information. Recent SOTA methods are mostly based on prompt learning, using lightweight prompts to encode degradation information. PromptIR [48] proposes a plug-and-play prompt module to guide the restoration process. DA-CLIP [36] utilizes a prompt learning module to incorporate degradation embeddings. MPerceiver [2] introduces a multi-modal prompt learning approach to harness Stable Diffusion priors. Despite achieving promising results, most existing methods use fixed network architectures, which may limit their adaptability to cover complex real-world scenarios.
Vision-Language Models. In recent years, vision-language models (VLMs) have shown strong performance across a wide range of multi-modal and vision-only tasks [51, 8, 29]. Among them, CLIP [51], as a powerful VLM, has demonstrated impressive zero-shot and few-shot capabilities across various high-level vision tasks [67, 93, 66]. However, in low-level vision tasks, CLIP’s capabilities have been relatively less explored. DA-CLIP [36] is the first to incorporate CLIP into all-in-one IR, employing a ControlNet-style [86] structure and using contrastive learning with image-text pairs to fine-tune CLIP. In this work, we focus on leveraging CLIP’s visual representation capabilities to efficiently capture degradation representations. Compared to DA-CLIP (Tab. 7), our proposed DG-Router requires 64×64$\times$64 × fewer learning parameters and 4×4$\times$4 × less training time, while achieving superior performance.
Parameter-efficient Fine-tuning. With the rise of large foundational models [51, 18, 1] in modern deep learning, the community has increasingly shifted its focus towards parameter-efficient fine-tuning (PEFT) methods for effective model adaptation. Among these, prompt learning [20, 90] and Low-Rank Adaptation (LoRA) [15] are two prominent and widely used PEFT methods. As discussed above, prompt learning has been widely applied in low-level vision tasks. LoRA posits that the weight changes during model adaptation follow a low-rank structure and incorporates trainable rank decomposition matrices into the pre-trained model. Specifically, the change matrix is re-parameterized into the product of two low-rank matrices: $W=W_{0}+\Delta W=W_{0}+sBA,$ where $W_{0}$ represents pre-trained weight matrix, $B\in\mathbb{R}^{m\times r}$ and $A\in\mathbb{R}^{r\times n}$ are low-rank matrices, $s=\frac{\alpha}{r}$ is the scaling factor. In this work, we first introduce LoRA into the all-in-one frameworks to facilitate efficient image restoration.
@@ 3 Method
As shown in Fig. 4, the image restoration network is based on the commonly used U-Net [65, 78, 4] structure, comprising stacked encoder, middle, and decoder blocks. LoRA-IR consists of two training stages: degradation-guided pre-training and parameter-efficient fine-tuning, both guided by the proposed Degradation-Guided Router (DG-Router). Following [3, 4], the model is optimized through PSNR loss. We first introduce the CLIP-based DG-Router in Sec. 3.1, which is used to extract robust degradation representations and provide probabilistic estimates to guide the training of LoRA-IR. Then we detail the pre-training process of LoRA-IR in Sec. 3.2. Finally, we describe the fine-tuning process in Sec. 3.3.
@@ 3.1 Degradation-guided Router
As shown in Fig. 4 (a), DG-Router uses a pre-trained CLIP image encoder to extract rich features from LQ images. The pre-trained CLIP image encoder typically limits input images to a smaller resolution (e.g., $224\times 224$). When handling higher-resolution images, a common approach [36] is to downsample the image to the resolution supported by CLIP using the processor. While this may have minimal impact on perception-based high-level classification tasks, significant downsampling can potentially lead to the loss of critical degradation information in pixel-level regression tasks like image restoration. Fig. 3 illustrates the results after the CLIP processor has processed the LQ images. Significant downsampling causes a substantial loss of degradation information, making it difficult to effectively extract degradation representations from the CLIP output features.
To address this issue, we propose a simple yet effective mechanism for scaling up the input resolution. For input LQ image $I_{LQ}\in\mathbb{R}^{H\times W\times 3}$, we use a sliding window to partition the image into small local patches $I_{slide}\in\mathbb{R}^{M\times H_{c}\times W_{c}\times 3}$, where $M$ is the number of patches, $H_{c}\times W_{c}$ denotes the resolution supported by CLIP. Both $I_{slide}$ and down-sampled image $I_{down}\in\mathbb{R}^{H_{c}\times W_{c}\times 3}$ are fed into the image encoder simultaneously, obtaining output features $e^{slide}\in\mathbb{R}^{M\times C_{clip}}$ and $e^{down}\in\mathbb{R}^{C_{clip}}$. As depicted in Fig. 4 (a), after pooling $e^{slide}$, we concatenate the features and pass them through a two-layer MLP to obtain the CLIP-extracted degradation embedding $e^{clip}$, which can be formulated as
@@
After feeding $e^{clip}$ into the classification head, we obtain the degradation prediction probabilities $w\in\mathbb{R}^{n}$, where $n$ is the number of degradation types. Without bells and whistles, the DG-Router is optimized using standard cross-entropy loss, with the only parameters being the classification head and the two-layer MLP. Once training is complete, all parameters of the DG-Router are frozen and no longer updated.
@@ 3.2 Degradation-guided Pre-training
where $\odot$ denotes the channel-wise multiplication, $\mathrm{MLP_{shared}}$ denotes the MLP projector shared across different blocks, $\mathrm{LN}$ denotes LayerNorm, $x_{in}$ is the original feature in the IR network, and $x_{out}$ is the feature after modulation. Through DAM modulation, the robust degradation representations from the DG-Router effectively enhance the degradation-specific knowledge of the IR network during pre-training.
Table 1: [Setting I] Quantitative comparisons for 4-task adverse weather removal. LoRA-IR surpasses recent SOTA all-in-one techniques, including MPerceiver [2] and Histoformer [56], across all evaluated datasets and metrics.
@@
LoRA-IR 33.39 0.9489
3.3 Parameter-efficient Fine-tuning
In the fine-tuning stage, we aim to utilize the Low-Rank Adaptation (LoRA) technique to model degradation characteristics and correlations efficiently, enhancing the model’s adaptability to real-world training-unseen degradations.
As shown in Fig. 4 (c), built upon the Mixture-of-Experts (MoE) architecture, we construct a set of low-rank restoration experts. We have a total of $n$ low-rank experts $\{E_{1},E_{2},\cdots,E_{n}\}$, where each expert is a learnable lightweight LoRA weight from the pre-trained restoration network in the first stage, specialized in handling a specific degradation type.
For a given input LQ image, the DG-Router predicts the degradation probability $w\in\mathbb{R}^{n}$, which serves as the score for selecting the appropriate experts for the restoration process. We sparsely select the top-$k$ highest-scoring experts as the most relevant ones, and achieve the final restoration result through their dynamic collaboration, formulated as
where $PreMod$ denotes the pre-trained module in the first stage, $\varphi(i)$ denotes the index of the $i$-th selected expert, $w^{\prime}\in\mathbb{R}^{n}$ represents the result of reapplying softmax normalization to the scores of the selected top-$k$ experts (with the weights of the unselected experts set to 00).
Note that the sparse selection mechanism in Eq. (3) grants LoRA-IR a self-adaptive network structure, enhancing its capacity to represent degradation-specific knowledge. The dynamic combination mechanism, on the other hand, enables collaboration among different restoration experts, effectively capturing the commonalities and correlations across various degradations. The design of the low-rank experts ensures the efficiency of LoRA-IR, allowing it to achieve high-performance all-in-one IR in a computationally efficient manner.
4 Experiments